The distance from a star to a planet is 8.1 * 10^17 m. How long does it take light, traveling at 10^16 m/year, to travel from the star to the planet?
Answer by checkley71(8403)

8.1*10^17/10^16=8.1*10^(17-16)=8.1*10 or 81 years.